OpenStudy (anonymous):

Find the first TWO non-negative asymptotes and the first negative asymptote of the graph of y in radians y=sec(5x). smallest non-negative asymptote: x= second non-negative asymptote: x= first negative asymptote: x= Find the first TWO non-negative asymptotes and the first negative asymptote of the graph of y in radians y=sec(5x). smallest non-negative asymptote: x= second non-negative asymptote: x= first negative asymptote: x=

OpenStudy (anonymous):

smallest non-negative asymptote: x= pi/6 second non-negative asymptote: x= pi/6 + pi/3 = pi/2 first negative asymptote: x= -pi/6 inflection point (on the midline at x=0 period = pi/3